### v2

# A Sobel-Gradient MLP Baseline for Handwritten Character Recognition

Abstract
We revisit the classical Sobel operator to ask a simple question: Are first-order edge maps sufficient to drive an all-dense MLP for handwritten character recognition(HCR), as an alternative to convolutional neural networks (CNNs)? Using only horizontal and vertical Sobel derivatives as input, we train a multilayer perceptron (MLP) on MNIST and EMNIST Letters. Despite its extreme simplicity, the resulting network reaches 98 % accuracy on MNIST digits and 92 % on EMNIST letters—approaching CNNs while offering a smaller memory footprint and transparent features. Our findings highlight that much of the class-discriminative information in handwritten character images is already captured by first-order gradients, making edge-aware MLPs a compelling option for HCR.
Keywords: handwritten recognition; Sobel operator; edge detection; multilayer perceptron; MNIST; EMNIST
1 Introduction
Handwritten-character recognition (HCR) is a long-standing testbed for vision, powering workflows from archival digitisation to on-device note transcription. Convolutional neural networks (CNNs) learn features directly from pixel intensities and achieve excellent accuracy. For digit and letter images, however, much of the class identity is determined by where stroke intensities change (edges) and in which direction they vary. This observation motivates revisiting explicit edge maps as inputs to a simple multilayer perceptron (MLP) classifier.
Why revisit Sobel now?
The Sobel–Feldman operator applies two fixed $3\times 3$ filters that approximate horizontal and vertical derivatives while providing mild smoothing. These maps highlight stroke contours with hardware-friendly computation and human-interpretable semantics. If a purely dense MLP can convert these fixed derivatives into decisions, the result is a model that is both lightweight and interpretable, providing a viable alternative to CNNs.
Research question.
Can a simple MLP, fed only with Sobel derivatives, approach standard CNN baselines on HCR datasets?
Contributions.
  1. 1.
    An MLP classifier that consumes only Sobel gradients and yields strong accuracy on MNIST and EMNIST Letters.
  2. 2.
    A minimal, reproducible pipeline (fixed preprocessing, fixed architecture, no augmentation) isolating the specific value of first-order edges.
  3. 3.
    Reproducible code and training protocol to facilitate follow-up work (see Code Availability).
2 Related Work
Gradient descriptors for HCR.
Before deep learning, edge descriptors such as Sobel/HOG paired with SVMs or $k$-NN achieved $95\%$ on MNIST [3, 8].
Edge-aware CNNs.
Several works prepend learnable edge filters to CNNs to bias features toward contours [11]. These remain convolutional and scale with feature-map depth; by contrast our parameters are fixed once the input dimension is set.
All-MLP architectures.
Token-mixing MLPs (e.g., MLP-Mixer [10]) show dense models can be competitive when inputs are suitably structured. We push this to an extreme: only two derivative channels. Relatedly, curvature-based descriptors have also been explored [4].
3 Background: The Sobel Gradient
3.1 Intuitive Explanation: Why Sobel Detects Edges
Think of an image as a checkerboard of tiny light-bulbs. The Sobel operator slides a $3\times 3$ window over this board and asks two simple questions at every position:
  1. 1.
    Left to right: Do the bulbs get noticeably brighter or darker as we move across?
  2. 2.
    Top to bottom: Do they get brighter or darker as we move downward?
If the answer to either question is “yes,” the centre pixel is on an edge. The two Sobel filters are hard-coded patterns that act like tiny digital feelers: one “feels” horizontal changes, the other vertical changes. Running them across the whole image produces two new maps—one for left-right edges, one for up-down edges—highlighting the outlines of strokes.
Why does this matter for a handwritten glyphs? Because the strokes are only a pixel or two thick; almost all the information that distinguishes a C from a G lies in these sharp intensity changes. By feeding the raw Sobel maps to an MLP, we give the network exactly those edge patterns without any extra noise or texture.
3.2 Discrete $3\times 3$ Kernels
The Sobel–Feldman operator [7] approximates the spatial image gradient by convolving the input $I$ with two separable kernels:
$G_{x}=\!\begin{bmatrix}-1&0&1\\[2.0pt] -2&0&2\\[2.0pt] -1&0&1\end{bmatrix},\qquad G_{y}=\!\begin{bmatrix}-1&-2&-1\\[2.0pt] 0&0&0\\[2.0pt] 1&2&1\end{bmatrix}$.
Both filters perform a central difference in one direction while embedding a mild Gaussian smoothing (weights 1:2:11\!:\!2\!:\!1) along the orthogonal axis. All coefficients are integers, which simplifies fixed-point or integer-only hardware implementations.
3.3 Interpreting the Output
Convolving $I$ with $G_{x}$ and $G_{y}$ yields horizontal and vertical derivative maps, respectively:
$\partial_{x}I=G_{x}*I,\qquad\partial_{y}I=G_{y}*I,$
where * denotes 2-D convolution. From these first-order derivatives one can derive the gradient magnitude $\lVert\nabla I\rVert=\sqrt{(\partial_{x}I)^{2}+(\partial_{y}I)^{2}}$ and the gradient orientation $\theta=\operatorname{atan2}(\partial_{y}I,\partial_{x}I)$. In this study we keep the raw signed derivatives $(\partial_{x}I,\partial_{y}I)$ because they preserve stroke polarity, which is useful for thin glyphs.
4 Method
Our goal is to isolate the value of first-order edges with a minimal pipeline: fixed derivatives $\to$ flatten $\to$ MLP.
Data flow and preprocessing
Images are $28\times 28$ grayscale, pixels are scaled to $[0,1]$, then passed to the Sobel operator to obtain $G_{x},G_{y}\in\mathbb{R}^{28\times 28}$ per image. Each channel is min–max normalised independently (min/max taken over that image and channel):
$\widehat{G}_{x}\;=\;\frac{G_{x}-\min(G_{x})}{\max(G_{x})-\min(G_{x})+\varepsilon},\qquad\widehat{G}_{y}\;=\;\frac{G_{y}-\min(G_{y})}{\max(G_{y})-\min(G_{y})+\varepsilon},\quad\varepsilon=10^{-8}.$
We stack $\widehat{G}_{x},\widehat{G}_{y}]\in\mathbb{R}^{2\times 28\times 28}$ and flatten to $\mathbf{x}\in\mathbb{R}^{1568}$ for the MLP.
Classifier
We use a three-layer MLP with batch normalisation (BN), ReLU, and dropout. The output dimension matches the task (10 for MNIST; 26 for EMNIST Letters):
Layer Dimension Components
Input 1568
Hidden 1 1024 Dense + BN + ReLU + Dropout(0.5)
Hidden 2 512 Dense + BN + ReLU + Dropout(0.4)
Hidden 3 256 Dense + BN + ReLU + Dropout(0.3)
Output (MNIST) 10 Dense + Softmax
Output (Letters) 26 Dense + Softmax
Parameter count (exact).
Including Dense biases and BN $\gamma,\beta$, the trainable parameters are:
[…]
Optimisation
  • Framework: TensorFlow/Keras 2.x
  • Optimiser: Adam (Keras defaults)
  • Loss function: sparse_categorical_crossentropy
  • Batch size: 128
  • Epochs: up to 50
  • Callbacks:
    • EarlyStopping on val_accuracy, patience = 4, restore_best_weights = True
    • ReduceLROnPlateau on val_loss, factor = 0.5, patience = 3, $\texttt{min\_lr}=10^{-6}$
  • Regularisation: No explicit L2 weight decay
  • Data augmentation: None
5 Experimental Setup
Datasets. MNIST digits [2] (70,000 images) and EMNIST Letters [1] (145,600 uppercase letters) are used. Data are loaded via TensorFlow Datasets (TFDS) [9].
Split protocol.
We concatenate TFDS train and test and perform a stratified 80/20 split using sklearn.model_selection.train_test_split. During training, Keras holds out 10% of the training portion as validation via $\texttt{validation\_split}=0.1$. Table 1 shows the resulting sizes.
Table 1: Stratified 80/20 split of TFDS train+test (with 10% of the training portion used for validation at fit time).
Dataset Total Train (fit) Val (held-out) Test
MNIST (0–9) 70,000 50,400 5,600 14,000
EMNIST Letters (A–Z) 145,600 104,832 11,648 29,120
Label handling.
For EMNIST Letters, TFDS yields labels 1–26; the implementation shifts to 0–25 by subtracting 1.
Evaluation metrics and class balance.
Both MNIST and the EMNIST Letters split used here have uniform per-class counts by design, and our 80/20 partition is stratified (§5), which preserves label frequencies. Accordingly, we report top-1 accuracy as the primary metric. If class imbalance were material in a deployment setting, we would additionally report macro-averaged F1, balanced accuracy, and per-class precision/recall; we omit these here due to the balanced label distribution.
Environment.
OpenCV for Sobel ($\texttt{cv2.Sobel}(\cdot,\cdot,1,0,\texttt{ksize}=3)$ and $\texttt{cv2.Sobel}(\cdot,\cdot,0,1,\texttt{ksize}=3)$); TensorFlow/Keras 2.x; TFDS; NumPy; scikit-learn.
6 Results
Table 2: Single-run test accuracy (%, rounded up) under the split protocol in Section 5.
Model MNIST EMNIST Letters
Sobel-Gradient MLP (this work) 98.0 92.0
Qualitative errors.
Errors cluster among visually similar letters with small stroke differences (e.g., C/G, I/J). The pattern is consistent with the intuition that first-order edges encode outline geometry well but may be ambiguous when final hooks or serifs decide the class.
7 Discussion
First-order edges capture core glyph structure.
The strong performance confirms that stroke contours alone encode sufficient cues for digit and letter classification.
Rotation and alignment.
The model is not rotation-invariant; modest rotations and off-centering degrade performance. Augmentations or rotation-normalised descriptors could address this.
Interpretability.
Inputs themselves are gradients, so introspection can directly highlight which signed-edge pixels most influence a decision (e.g., mapping large-magnitude input dimensions or first-layer weights back to image space) without a separate saliency pipeline [6].
Limitations.
(i) No augmentation or robustness measurements (noise/blur/thickness) are reported. (ii) No latency/energy measurements are reported here; on-device claims should be backed by measured timings. (iii) We do not evaluate larger alphabets (e.g., EMNIST Balanced) or scripts beyond Roman uppercase; generalisation remains to be shown.
8 Conclusion
A compact MLP fed only with Sobel derivatives achieves strong accuracy on MNIST and EMNIST Letters under a simple training recipe. The approach offers transparent, geometry-aligned inputs and a predictable parameter footprint, providing a clear baseline for studying how much local edge geometry a recogniser really needs.
Future work.
Add controlled robustness ablations (rotation, noise, blur, thickness), study derivative encodings (signed vs. magnitude/orientation), extend to larger benchmarks (EMNIST Balanced, NIST SD19), and report on-device latency/energy with quantisation.
Code Availability
All code and training scripts are available at
https://github.com/AzamNouri/Sobel-Gradient-MLP.
For reproducibility, an archived release is preserved on Zenodo [5].